# EditGRPO: Reinforcement Learning with Post‑Rollout Edits for Clinically Accurate Chest X‑Ray Report Generation

### v1

@@ Appendix E Training Configurations
In the reinforcement learning setup, we use a rollout batch size of 32. Image inputs are constrained to a maximum of 401,408 pixels and a minimum of 262,144 pixels. We employ the AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1.0 e-6 and a weight decay of 1.0 e-2. During the rollout phase, the generation temperature is set to 1.0, and we sample $\frac{G}{2}=5$ responses per prompt. To reduce memory usage, we enable gradient checkpointing (Chen et al., 2016) and apply Fully Sharded Data Parallelism (FSDP) (Zhao et al., 2023). Additionally, vLLM (Kwon et al., 2023) is used to accelerate inference and response generation. We set the cosine similarity threshold $\tau$ to 0.6 in the editing rules based on preliminary observations, without formal hyperparameter optimization.
@@ Appendix F Datasets
Our study employs four fully de‑identified, publicly released chest radiograph collections—MIMIC‑CXR (Johnson et al., 2019), PadChest‑GR (de Castro et al., 2025) (English version), IU‑Xray (Demner-Fushman et al., 2016), and RexGradient (Zhang et al., 2025a)—each governed by permissive open‐use terms that together guarantee ethical compliance and reproducibility. The MIMIC‑CXR dataset is accessed under its data use agreement and CITI “Data or Specimens Only Research” certification. PadChest‑GR is distributed under the PADCHEST Dataset Research Use Agreement [^5] https://bimcv.cipf.es/bimcv-projects/padchest/padchest-dataset-research-use-agreement/, which grants free access for academic research only and explicitly prohibits redistribution or commercial use. IU‑Xray is released under a CC BY‑NC‑ND 4.0 license, which permits non‑commercial redistribution with attribution but prohibits both commercial reuse and the creation of derivative works. RexGradient is available via Hugging Face [^6] https://huggingface.co/datasets/rajpurkarlab/ReXGradient-160K, and is released under a Non‑Commercial Data Access and Use Agreement that restricts use to non‑clinical, non‑commercial research while requiring proper attribution to Harvard Medical School and Gradient Health. No patient can be re‑identified from any of these sources, and all secondary analyses conform strictly to each dataset’s specific attribution, share‑alike, and noncommercial clauses. By restricting our work to these four open‑license datasets, we uphold the highest standards of data privacy, transparency, and legal clarity in ethical AI research.
Appendix G Additional Experiments
In this section, we present comprehensive evaluation metrics on IU-XRay and PadChest-GR, as shown in Table 5. In addition, we conduct ablation studies on both reward design (see Fig. 3) and post-rollout editing strategies (see Fig. 4) to analyze their individual contributions to overall performance.
Figure 3: Influence of reward design on the IU-XRay dataset under the SFT + Dr.GRPO setting. RG denotes the RadGraph reward, and IF denotes the inverse‑frequency reward, which assigns higher scores when a rare condition is hit according to the label distribution across the 14 CheXpert classes (Irvin et al., 2019) of the training data.
Composite reward design makes sense.
@@ Appendix G Additional Experiments
Although RaTEScore only identifies presence or absence (as abnormality versus non-abnormality or disease versus non-disease), RadGraph identifies three presence labels (definitely present, definitely absent, or uncertain). However, RadGraph does not come with an entity embedding model, so the cosine similarity threshold reverts to exact entity matching. Results on IU-XRay, shown in Table 4, indicate that RadGraph-based editing underperforms. This likely stems from reliance on exact string matching with RadGraph, which lacks canonical concept IDs or mention‑level embeddings for semantic matching, making the procedure sensitive to phrasing (Jain et al., 2021).

### v2

@@ Appendix D Training Configurations
In the reinforcement learning setup, we use a rollout batch size of 32. Image inputs are constrained to a maximum of 401,408 pixels and a minimum of 262,144 pixels. We employ the AdamW (loshchilov2017decoupled) optimizer with a learning rate of 1.0 e-6 and a weight decay of 1.0 e-2. During the rollout phase, the generation temperature is set to 1.0, and we sample $\frac{G}{2}=5$ responses per prompt. To reduce memory usage, we enable gradient checkpointing (chen2016training) and apply Fully Sharded Data Parallelism (FSDP) (zhao2023pytorch). Additionally, vLLM (kwon2023efficient) is used to accelerate inference and response generation.
@@ Appendix E Datasets
Our study employs four fully de‑identified, publicly released chest radiograph collections—MIMIC‑CXR (mimiccxr; johnson2019mimic), PadChest‑GR (de2025padchest) (English version), IU‑Xray (iuxray), and RexGradient (zhang2025rexgradient)—each governed by permissive open‐use terms that together guarantee ethical compliance and reproducibility. The MIMIC‑CXR dataset is accessed via PhysioNet under its data use agreement and CITI “Data or Specimens Only Research” certification goldberger2000physiobank. PadChest‑GR is distributed under the PADCHEST Dataset Research Use Agreement [^5] https://bimcv.cipf.es/bimcv-projects/padchest/padchest-dataset-research-use-agreement/, which grants free access for academic research only and explicitly prohibits redistribution or commercial use. IU‑Xray is released under a CC BY‑NC‑ND 4.0 license, which permits non‑commercial redistribution with attribution but prohibits both commercial reuse and the creation of derivative works. RexGradient is available via Hugging Face [^6] https://huggingface.co/datasets/rajpurkarlab/ReXGradient-160K, and is released under a Non‑Commercial Data Access and Use Agreement that restricts use to non‑clinical, non‑commercial research while requiring proper attribution to Harvard Medical School and Gradient Health. No patient can be re‑identified from any of these sources, and all secondary analyses conform strictly to each dataset’s specific attribution, share‑alike, and noncommercial clauses. By restricting our work to these four open‑license datasets, we uphold the highest standards of data privacy, transparency, and legal clarity in ethical AI research.
Appendix F Additional Analysis
In this section, we present comprehensive evaluation metrics on IU-XRay and PadChest-GR, as shown in Table 5. In addition, we conduct ablation studies on both reward design (see Fig. 3) and post-rollout editing strategies (see Fig. 4) to analyze their individual contributions to overall performance. Detailed results of the statistical significance tests are presented in Table 8.
Figure 3: Influence of reward design on the IU-XRay dataset under the SFT + Dr.GRPO setting. RG denotes the RadGraph reward, and IF denotes the inverse‑frequency reward, which assigns higher scores when a rare condition is hit according to the label distribution across the 14 CheXpert classes (chexpert) of the training data.
Composite reward is effective.
@@ Appendix F Additional Analysis
Although RaTEScore only identifies presence or absence (as abnormality versus non-abnormality or disease versus non-disease), RadGraph identifies three presence labels (definitely present, definitely absent, or uncertain). In our implementation, instances marked uncertain are excluded. However, RadGraph does not come with an entity embedding model, so the cosine similarity threshold reverts to exact entity matching. Results on IU-XRay, shown in Table 4, indicate that RadGraph-based editing underperforms. This likely stems from reliance on exact string matching with RadGraph, which lacks canonical concept IDs or mention‑level embeddings for semantic matching, making the procedure sensitive to phrasing (radgraph).
Similarity threshold selection.
We set the cosine similarity threshold $\tau$ in the editing rules from the preliminary ablation study using a 3,000-case subset of the MIMIC-CXR, with evaluation on the full test set. As shown in the Table 7, $\tau$=0.6 yields the best performance across five of the seven metrics. CheXbert-14-Micro and Macro scores, which are directly associated with the model’s ability to detect common abnormalities, show the most significant improvements. This suggests that $\tau$=0.6 effectively balances the trade-off between different editing operations. For the RaTE and RadGraph metrics, we observed that $\tau$=0.9 achieved the best performance, although the improvement was marginal. This is consistent with the nature of these metrics, particularly RadGraph, which is more sensitive to exact wording and literal entity agreement. A higher $\tau$ enforces a stricter entity-matching criterion, which slightly benefits these metrics. However, our results confirm that a balanced, sentence-level editing approach enabled by a moderate $\tau$ is crucial for maximizing overall clinical efficacy across a range of complementary metrics, as demonstrated by the superior performance on CheXbert and GREEN scores.
Metrics $\tau=0$ $\tau=0.3$ $\tau=0.6$ $\tau=0.9$
RaTEScore 0.5290 0.5333 0.5324 0.5338
RadGraph F1 0.2720 0.2745 0.2726 0.2769
Micro-F1-14 0.5071 0.4995 0.5124 0.5032
Macro-F1-14 0.3374 0.3290 0.3531 0.3299
Micro-F1-5 0.5477 0.5422 0.5533 0.5482
Macro-F1-5 0.4656 0.4569 0.4712 0.4707
GREEN 0.3389 0.3368 0.3455 0.3433
Table 7: Influence of $\tau$ on MIMIC-CXR subset.
Table 8: Per-metric deltas of EditGRPO against SFT and Dr.GRPO across three datasets. Cells show absolute differences; marks per-metric significance vs SFT and vs Dr.GRPO (two-sided tests as defined in the text). The Avg. row is excluded.
MIMIC-CXR IU-XRAY PadChest-GR
Metric $\Delta$ Edit–SFT $\Delta$ Edit–Dr $\Delta$ Edit–SFT $\Delta$ Edit–Dr $\Delta$ Edit–SFT $\Delta$ Edit–Dr
RaTE +0.0278 +0.0252 +0.0226 +0.0164 +0.1100 +0.0769
CheXbert_14_Micro +0.0450 +0.0124 -0.0130 -0.0046 +0.0075 +0.0480
CheXbert_14_Macro +0.0330 +0.0193 +0.0165 +0.0049 +0.0385 -0.0041
CheXbert_5_Micro +0.0499 +0.0074 +0.0839 +0.0696 +0.0903 +0.0065
CheXbert_5_Macro +0.0505 +0.0219 +0.0502 +0.0449 +0.0652 +0.0149
GREEN +0.0242 +0.0198 +0.0711 +0.0535 +0.0066 +0.0132
RadGraph F1 -0.0096 -0.0118 +0.0457 +0.0149 -0.0005 +0.0393
Statistical significance testing.
We assess per-metric differences $\Delta(\text{Edit--SFT})$ and $\Delta(\text{Edit--Dr.GRPO})$ using two-sided paired tests at the sample level; Table 8 reports absolute deltas and marks significance with (vs SFT) and (vs Dr.GRPO) for the seven primary metrics (RaTE, CheXbert-14 Micro/Macro, CheXbert-5 Micro/Macro, GREEN, RadGraph F1; Avg. excluded). On MIMIC-CXR, EditGRPO is significant on 6/7 metrics vs SFT and 5/7 vs Dr.GRPO; on IU-XRAY, 6/7 vs SFT and 6/7 vs Dr.GRPO; on PadChest-GR, 5/7 vs SFT and 6/7 vs Dr.GRPO. The largest and most frequent gains occur on CheXbert-5 (micro/macro) and GREEN, while RadGraph F1 shows a small non-significant drop on MIMIC-CXR and reaches significance only vs Dr.GRPO on PadChest-GR, consistent with its sensitivity to relational structure.